# Single Bell inequality to detect genuine nonlocality in three-qubit genuinely entangled states

### v1

@@ Abstract
It remains an open question whether every pure multipartite state that is genuinely entangled is also genuinely nonlocal. Recently, a new general construction of Bell inequalities allowing the detection of genuine multipartite nonlocality (GMNL) in quantum states was proposed in [F. J. Curchod, M. L. Almeida, and A. Acín, New J. Phys. 21, 023016 (2019)] with the aim of addressing the above problem. Here we show how, in a simple manner, one can improve this construction to deliver tighter Bell inequalities for detection of GMNL. Remarkably, we then prove one of the improved Bell inequalities to be powerful enough to detect GMNL in every three-qubit genuinely entangled state. We also generalize some of these inequalities to detect not only GMNL but also nonlocality depth in multipartite states and we present a possible way of generalizing them to the case of more outcomes.
@@ I Introduction
Our aim in this work is to show that a slightly improved version of one of the inequalities of Ref. [21] is powerful enough to definitely settle the problem of equivalence of genuine nonlocality and genuine entanglement for three-qubit states. In other words, we introduce a single Bell inequality that is capable of detecting genuine nonlocality in any three-qubit genuinely entangled. We thus not only recover and confirm the result of Ref. [15], but we also introduce a universal test of genuine nonlocality for three-qubit states which, unlike the Hardy-type arguments, is capable of tolerating some amount of experimental imperfections and noises. On the way of achieving the above result we show how to modify the construction of Ref. [21] to provide tighter Bell inequalities detecting GMNL which improves that of Ref. [21]. We also generalize some of the inequalities of Ref. [21] to detect the nonlocality depth of multipartite systems. Finally, we explore the possibility of generalizing the approach of Ref. [21] to Bell scenarios involving more outcomes, concentrating on a particular case of three-outcome measurements. Indeed, we show how to suitably modify the CGLMP Bell inequality [24] to derive an inequality detecting GMNL in tripartite systems which is analogous to inequalities of Ref. [21]
@@ II.1 Genuinely multipartite entanglement and nonlocality
$|\mathrm{GHZ}_{n,d}\rangle=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle^{\otimes n}.$ (1)
@@
where $r_{S|\bar{S}}$ is some probability distribution.
@@
$p(\boldsymbol{a}|\boldsymbol{x})=\sum_{\lambda}q_{\lambda}^{S|\bar{S}}p_{S}(\boldsymbol{a}_{S}|\boldsymbol{x}_{S},\lambda)p_{\bar{S}}(\boldsymbol{a}_{\bar{S}}|\boldsymbol{x}_{\bar{S}},\lambda),$ (6)
@@ II.2 Class of Bell inequalities for detection of GMNL
$I^{i,j}_{\boldsymbol{0}|\boldsymbol{0}}=$ $p(\boldsymbol{0}|\boldsymbol{0})-p(0\ldots 01_{i}0\ldots 0|0\ldots 01_{i}0\ldots 0)$ (11)
@@
It is worth mentioning that, as indicated in Ref. [21], any other, in general $m$-partite Bell inequality with $m<n$ in the simplest Bell scenario with two dichotomic measurements per party, which can be written as
@@
where $\beta_{\mathbf{x}}^{\mathbf{a}}\geqslant 0$, can be used as a seed inequality for the construction. In Ref. [21] a few examples of Bell inequalities in the bipartite and multipartite scenarios were presented which admit the above decomposition such as for instance the one found in Ref. [14] which reads
@@ III.1 General construction
    To formulate the second condition, we first decompose $\mathcal{I}^{g}$ as
@@
    where in both expressions ±$\mathcal{I}^{g}_{\pm}$ probabilities enter with positive weights. In other words, we decompose each Bell expression into two terms containing probabilities that enter the Bell expression with positive and negative weights, respectively. Then, the second condition on expressions $\mathcal{I}^{g}$ requires that the positive parts of every Bell expression $\mathcal{I}^{g}$ are the same, that is,
    $\forall_{g\in G}\qquad\mathcal{I}^{g}_{+}=\mathcal{I}_{+}.$ (19)
Let us also denote by $T$ the common terms appearing in all the negative parts $\mathcal{I}_{-}^{g}$. All together this implies that for all $g\in G$ we have
@@
Our aim now is to determine an upper bound for the expression (17) for bilocal models (7). For this purpose, we consider a particular bipartition $S|\bar{S}$. Then, for all $\vec{p}$ local across the bipartition $S|\bar{S}$ (6) we have
@@
where $G^{\prime}\subseteq G$ is a set of such $g$ that either $\{A_{i}\}_{i\in g}\subset S$ or $\{A_{i}\}_{i\in g}\subset\bar{S}$. Next, using the decomposition (19), and Ineq. (20), the above can further be upper bounded by
@@
Consider a Bell expression (17) such that the conditions (i) and (ii) are satisfied by $I^{g}$ for each $g\in G$. Then, violation of the following Bell inequality
@@
In order to demonstrate that the above inequality is tighter than that in (13) we consider a mixture of the $n$-qubit GHZ state (1) and white noise,
$\rho(q)=(1-q)|\mathrm{GHZ}_{n,2}\rangle\!\langle\mathrm{GHZ}_{n,2}|+q\frac{\mathds{1}}{2^{n}}$ (26)
with $q\in[0,1]$. The numerical comparison between noise levels needed for violation of either inequality has been presented in Fig. 1. The numerical analysis shows that, in the case of a three-qubit GHZ state, the violation of Ineq. (13) is achieved for noise levels of $q<0.05$ while for the new inequality (25) we can observe violation for $q<0.106$.
@@
The maximal violations of the inequalities were found assuming projective measurements. Each was parameterized by two real angles and then the standard methods in Mathematica were used to find the maxima.
@@ III.3 Detection of nonlocality depth
The task is to find the $L_{k}$ partition that maximizes the number of $I^{i,j}_{\boldsymbol{0}|\boldsymbol{0}}$ terms between parties from the same subset, which can be upper-bounded by $I_{\mathbf{0}|\mathbf{0}}^{i,j}\leqslant p(\mathbf{0}|\mathbf{0})$. The optimal strategy for maximizing this number is to take the partition for which $|S_{i}|=k$ for all $i\in\{1,\ldots,L-1\}$ and $|S_{L}|=n\mod k$, where $L=\lceil n/k\rceil$. We thus obtain the following inequality
@@ III.4 A possible generalization to more outcomes
$\mathcal{I}_{+}$ $=$ $\mathcal{I}_{+}^{AB}=\mathcal{I}_{+}^{AC}=\mathcal{I}_{+}^{BC}$
@@ IV Conclusions
There is a couple of ways in which the above research line can be continued. First, generalizing our proof for tripartite genuinely entangled states, the inequalities we introduced can perhaps be used to prove that all pure, genuinely entangled states are genuinely nonlocal which is a long-standing conjecture. Next, there is a question of how to maximally increase the common negative term giving a greater noise-robustness of the inequality. Here, a couple of possible approaches are possible. One could, for example, rewrite the ”seed” inequalities using non-signaling assumption, or perhaps lift these inequalities using different measurement settings and results, akin to what we did for $3$-outcome inequality (45).
@@ Appendix A Proof of Theorem 2
By the virtue of [27, Theorem 3], every pure, $3$-partite, qubit state can be written as
@@
Furthermore, in Ref. [21] it was shown that the inequality (13) is violated by all pure, genuinely multipartite entangled states that are invariant under the permutation of a pair of qubits. As the new inequality (25) is tighter than (13), it directly follows that it also has to be violated by such states. Therefore, we can restrict ourselves to states that are not invariant under the permutation of any two qubits, which implies that $b>c>d$.

### v3

# Single Bell inequality to detect genuine nonlocality in three-qubit pure genuinely entangled states

@@ Abstract
It remains an open question whether every pure multipartite state that is genuinely entangled is also genuinely nonlocal. Recently, a new general construction of Bell inequalities allowing the detection of genuine multipartite nonlocality (GMNL) in quantum states was proposed in [F. J. Curchod, M. L. Almeida, and A. Acín, New J. Phys. 21, 023016 (2019)] with the aim of addressing the above problem. Here we show how, in a simple manner, one can improve this construction to deliver finer Bell inequalities for detection of GMNL. Remarkably, we then prove one of the improved Bell inequalities to be powerful enough to detect GMNL in every three-qubit genuinely entangled state. We also generalize some of these inequalities to detect not only GMNL but also nonlocality depth in multipartite states and we present a possible way of generalizing them to the case of more outcomes.
@@ I Introduction
Our aim in this work is to show that a slightly improved version of one of the inequalities of Ref. [21] is powerful enough to definitely settle the problem of equivalence of genuine nonlocality and genuine entanglement for three-qubit states. In other words, we introduce a single Bell inequality that is capable of detecting genuine nonlocality in any three-qubit genuinely entangled. We thus not only recover and confirm the result of Ref. [15], but we also introduce a universal test of genuine nonlocality for three-qubit states which, unlike the Hardy-type arguments which require a certain set of equalities to be satisfied, is capable of tolerating some amount of experimental imperfections and noises. On the way of achieving the above result we show how to modify the construction of Ref. [21] to provide finer Bell inequalities detecting GMNL which improves that of Ref. [21]. We also generalize some of the inequalities of Ref. [21] to detect the nonlocality depth of multipartite systems. Finally, we explore the possibility of generalizing the approach of Ref. [21] to Bell scenarios involving more outcomes, concentrating on a particular case of three-outcome measurements. Indeed, we show how to suitably modify the CGLMP Bell inequality [24] to derive an inequality detecting GMNL in tripartite systems which is analogous to inequalities of Ref. [21]
@@ II.1 Genuinely multipartite entanglement and nonlocality
$|\mathrm{GHZ}_{n,d}\rangle:=\frac{1}{\sqrt{d}}\sum_{i=0}^{d-1}|i\rangle^{\otimes n}.$ (1)
@@
where $r_{S|\bar{S}}$ is some probability distribution and the sum is over all nontrivial bipartitions for which both sets $S$ and $\bar{S}$ are nonempty.
@@
$p(\boldsymbol{a}|\boldsymbol{x})=\sum_{\lambda}q_{\lambda}p_{S}(\boldsymbol{a}_{S}|\boldsymbol{x}_{S},\lambda)p_{\bar{S}}(\boldsymbol{a}_{\bar{S}}|\boldsymbol{x}_{\bar{S}},\lambda),$ (6)
@@ II.2 Class of Bell inequalities for detection of GMNL
$I^{i,j}_{\boldsymbol{0}|\boldsymbol{0}}:=$ $p(\boldsymbol{0}|\boldsymbol{0})-p(0\ldots 01_{i}0\ldots 0|0\ldots 01_{i}0\ldots 0)$ (11)
@@
It is worth mentioning that in the simplest Bell scenario involving two dichotomic measurements per observer the above procedure can be phrased in a more general way [21]. In fact, any $m$-partite Bell inequality with $m<n$ which admits the following form
@@
can be used as a seed inequality for the construction.
In Ref. [21] a few examples of Bell inequalities in the bipartite and multipartite scenarios were presented which admit the above decomposition such as for instance the one found in Ref. [14] which reads
@@ III.1 General construction
    To formulate the second condition, we first decompose $\mathcal{I}^{g}$ as a difference of two expressions,
@@
    where both ±$\mathcal{I}^{g}_{\pm}$ only contain probabilities with positive weights. In other words, we decompose each expression $\mathcal{I}^{g}$ into two expressions containing probabilities that enter the $\mathcal{I}^{g}$ with positive and negative signs, respectively. Then, the second condition on expressions $\mathcal{I}^{g}$ requires that the positive parts of every Bell expression $\mathcal{I}^{g}$ are the same, that is,
    $\forall_{g\neq g^{\prime}\in G}\qquad\mathcal{I}^{g}_{+}=\mathcal{I}^{g^{\prime}}_{+}=:\mathcal{I}_{+}.$ (19)
Let us also denote by $T$ the combination of probabilities that is common to all the negative parts $\mathcal{I}_{-}^{g}$. In other words, we decompose all $\mathcal{I}_{-}^{g}$ as a sum of $T$ and the remaining terms that are different for different $\mathcal{I}_{-}^{g}$; in particular, if $\mathcal{I}_{-}^{g}$ do not share any common terms, $T=0$. All together this implies that for all $g\in G$ we have
@@
Our aim now is to determine an upper bound for the expression (17) for bilocal models (7). For this purpose, we consider a particular bipartition $S|\bar{S}$. Then, for all $\vec{p}$ that are local across the bipartition $S|\bar{S}$ (6) we have
@@
where $G^{\prime}\subseteq G$ is defined as
$G^{\prime}=\{g\in G\,|\,\{A_{i}\}_{i\in g}\subset S\,\mathrm{or}\,\{A_{i}\}_{i\in g}\subset\bar{S}\};$ (22)
for the remaining $g\in G$ such that $g\notin G^{\prime}$ we employed the condition (i) stated above. Next, using the decomposition (19), and Ineq. (20), the above can further be upper bounded by
@@
Consider a Bell expression (17) such that the conditions (i) and (ii) are satisfied by $\mathcal{I}^{g}$ for each $g\in G$. Then, violation of the following Bell inequality
@@
In order to demonstrate that the above inequality is finer than that in (13) we consider a mixture of the $n$-qubit GHZ state (1) and white noise,
$\rho(q):=(1-q)|\mathrm{GHZ}_{n,2}\rangle\!\langle\mathrm{GHZ}_{n,2}|+q\frac{\mathds{1}}{2^{n}}$ (27)
with $q\in[0,1]$. The numerical comparison between noise levels needed for violation of either inequality is presented in Fig. 1. The numerical analysis shows that, in the case of a three-qubit GHZ state, the violation of Ineq. (13) is achieved for noise levels of $q<0.05$ while for the new inequality (26) we can observe violation for $q<0.106$.
@@
Each point in Fig. 1 was obtained by optimizing the Bell expressions (13) and (26) over two pairs of dichotomic observables, one pair corresponding to party $A_{1}$ and the other pair corresponding to the remaining parties; here we assume that all parties except $A_{1}$ measure the same pair of observables which is motivated by the fact that the noisy GHZ state $\rho(q)$ are permutationally invariant. Each observable was parameterized by two real angles and then the standard methods in Mathematica were used to find the maxima [27].
@@ III.3 Detection of nonlocality depth
The task is to find the $L_{k}$ partition that maximizes the number of $I^{i,j}_{\boldsymbol{0}|\boldsymbol{0}}$ terms between parties from the same subset, which can be upper-bounded by $I_{\mathbf{0}|\mathbf{0}}^{i,j}\leqslant p(\mathbf{0}|\mathbf{0})$. Let us consider an $L_{k}$ partition for which there exist two sets $S_{l_{1}},S_{l_{2}}$ such that $|S_{l_{1}}|\leqslant|S_{l_{2}}|<k$. It is easy to show that a different $L_{k}$ partition for which $S_{l_{2}}^{\prime}=S_{l_{2}}\cup\{s\}$, $S_{l_{1}}^{\prime}=S_{l_{1}}\setminus\{s\}$, where $s\in S_{l_{1}}$ and $S_{l}^{\prime}=S_{l}$ for any other $l$, gives a higher number of such $I^{i,j}_{\boldsymbol{0}|\boldsymbol{0}}$ terms. This implies that the optimal strategy for maximizing this number is to take the partition for which $|S_{i}|=k$ for all $i\in\{1,\ldots,L-1\}$ and $|S_{L}|=n\mod k$, where $L=\lceil n/k\rceil$. We thus obtain the following inequality
@@ III.4 A possible generalization to more outcomes
$\mathcal{I}_{+}$ ::=:= $\mathcal{I}_{+}^{AB}=\mathcal{I}_{+}^{AC}=\mathcal{I}_{+}^{BC}$
@@ IV Conclusions
There is a couple of ways in which the above research line can be continued. First, generalizing our proof for tripartite genuinely entangled states, the inequalities we introduced can perhaps be used to prove that all pure, genuinely entangled states are genuinely nonlocal which is a long-standing conjecture. Next, there is a question of how to maximally increase the common negative term giving a greater noise-robustness of the inequality. Here, a couple of approaches are possible. One could, for example, rewrite the ”seed” inequalities using non-signaling assumption, or perhaps lift these inequalities using different measurement settings and results, akin to what we did for $3$-outcome inequality (46). Another interesting question is how to generalize this approach to deliver Bell inequalities detecting GMNL in scenarios involving arbitrary number of outcomes and thus being applicable to multipartite states composed of many qudits. While we have made the first step towards achieving this goal by using the three-outcome CGLMP Bell inequality, how to obtain a general construction remains open. A related question is whether using the SATWAP Bell inequality can be advantageous in this direction [31].
@@
V Acknowledgments
We acknowledge the support by the (Polish) National Science Center through the SONATA BIS Grant No. 2019/34/E/ST2/00369.
@@ Appendix A Proof of Theorem 2
By the virtue of [28, Theorem 3], every pure, $3$-partite, qubit state can be written as
@@
[…]